OpenStudy (anonymous):

how do you change from cartesian coordinates to polar coordinates when you x=0, because you cant use tan=y/x

OpenStudy (gorv):

well cordinate means (x,y) we need both to convert

OpenStudy (gorv):

tan(pi/2)= sin(pi/2)/cos(pi/2)=1/0=infinity so not a probe tan theta =y/x=y/0=infinity means theta = pi/2

OpenStudy (gorv):

but for r u need to be known y value